For the function
f(x)=−3−5x^3

find the derivative of the inverse function f-1 at the point x=37.

(f−1)'(37)=
Answer #1

(00. Q. f(x) = -3- 5x3 에 3 4 - - 3-53 -3 = 3+ sx 2 5X3 3해 N * Sx? = -3-y 83 -3-J 음 외2 N * (3) 크성에 이루지 (fai] + 글(SS}ty (f(x))